OpenStudy (vuriffy):

Have a test tomorrow, worksheet has a few questions we haven't gone over and I know my teacher will put them on there, because she always does that:

OpenStudy (vuriffy):

\[\log_{10}(x+9)=2+\log_{10}x \]

OpenStudy (vuriffy):

Solve.

OpenStudy (vuriffy):

I am still unsure on how to start it.

Hayhayz (hayhayz):

just subtract

Hayhayz (hayhayz):

log_10(x+9)-log_10(x) =2 so log_10((x+9)/x) = 2 (x+9)/x = 10^2 x+9=100x x=9/99 =1/11

OpenStudy (vuriffy):

So, you turn it around to equal the whole number?

Hayhayz (hayhayz):

remmber that log(x) + log(y) = log(x*y) and log(x)-log(y)=log(x/y) the basic log rules

Hayhayz (hayhayz):

you can try to prove these identities to yourself, it will make you more comfortable with it

OpenStudy (vuriffy):

I understand it now, thank you. I just didn't know how to start it. I have two more, if you don't mind helping me with them. I can tag you.

Hayhayz (hayhayz):

sure :)